OpenStudy (anonymous):

A rectangle has a length of n units. Its width is 8 units more than its length. What is the area of the rectangle?

OpenStudy (anonymous):

Length = n Width = length + 8 = n + 8 Area = length (width) = n(n + 8) = n^2 + 8n units^2

OpenStudy (anonymous):

Area of a rectangle = length * width = n * (n + 8) =n^2 + 8n units

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Thank you so much